Using the appropriate interest table, compute the present values of the following periodic amounts due at the end of the designated periods.

$54,230 receivable at the end of each period for 7 periods compounded at 12%.

$54,230 payments to be made at the end of each period for 17 periods at 9%. (

$54,230 payable at the end of the seventh, eighth, ninth, and tenth periods at 12%.
Answer #1

a) Present value of an annuity of 1 for 7 periods @ 12% = 4.5638

Periodic receivables = 54230

Present value of receivables = 54230*4.5638 = 247494.87

b )Present value of an annuity due of 1 for 17 periods at 9% = 9.3126

Periodic payments = 54230

Present value of payments = 54230 * 9.3126 = 505022.30

c) Present value of an annuity due of 1 for 6 periods at 12% = 4.1114

Present value of an annuity due of 1 for 10 periods at 12% = 5.6502

Periodic payments = 54230

Present value of payments = 54230*(5.6502-4.1114) = 83449.12/-